# MGPATH: Vision-Language Model with Multi-Granular Prompt Learning for Few-Shot WSI Classification

### v1

@@ 3.3.1 Patches-based Prompting
The frozen image encoder $E_{I}(.)$ (or $A_{I}(E_{I}(.))$ in case of GigaPath-PLIP) is used to map patches $I$ into a feature vector $H=\{H^{(l)}\in\mathbb{R}^{N_{l}\times d},H^{(h)}\in\mathbb{R}^{N_{h}\times d}\}$ where $d$ denotes the feature dimension. To effectively consolidate the extensive set of patch features into a final slide-level representation, we introduce a set of learnable visual prompts $\mathbf{p}_{v}\in\mathbb{R}^{N_{p}\times d}$, which facilitate the progressive merging of patch features in $H^{(l)}$(similarly for $H^{(h)}$) (Figure 2). In particular, we formulate $\mathbf{p}_{v}$ as Query and take all features in $H^{(l)}$ as the Keys $K^{(l)}_{p}$ and Values $V^{(l)}_{p}$ in in self-attention [62]. We then associate $\mathbf{p}_{v}$ with patch features as:
@@ 3.3.2 Spatial Patch Group-based Prompting
We next design a trainable message-passing network $g_{\epsilon}(.)$ based on the graph attention layer (GAT) [63] to capture the feature representation of each node and its local neighbors. The message passing of the GAT layer is formulated as:
@@ 3.4 Optimal Transport for Visual-Text Alignment
Given descriptive text prompts $\mathbf{T}^{(l)}$ and $\mathbf{T}^{(h)}$ (Eq.(2)) and visual prompt-guided slide features $\mathbf{p}_{v}^{(l)}$ and $\mathbf{p}_{v}^{(h)}$ (Eq.(6)) for low and high resolutions, our goal is to maximize the similarity between slide and text embeddings for each class $c$. Rather than relying on cosine distance, as in prior works [75, 74, 72, 46, 46, 56], we propose using optimal transport (OT)-based distance to capture a more nuanced cross-alignment between visual and text domains. Although OT has been explored for prompt learning in natural images and multi-modal learning [24, 5, 38, 51], we are the first to adapt it for whole-slide imaging (WSI), effectively handling the alignment of multi-magnification patches to capture rich structural details across scales.
@@ 4.1 Settings
PatchGastricADC22[61] consists of approximately 262K patches derived from WSI of H&E-stained gastric adenocarcinoma specimens, each paired with associated diagnostic captions collected from the University of Health and Welfare, Mita Hospital, Japan. QUILT-1M [20] includes approximately 653K images and one million pathology image-text pairs, gathered from 1,087 hours of educational histopathology videos presented by pathologists on YouTube. ARCH [15] is a pathology multiple-instance captioning dataset containing pathology images at the bag and tile level. However, our work focuses on tile-level images from all datasets for our contrastive training strategy. In total, we collected approximately 923K images from these datasets.
@@
For the classification task, the proposed method was evaluated in three datasets from the Cancer Genome Atlas Data Portal[60]: TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA. We followed the ViLa-MIL[53] experimental settings for TCGA-NSCLC and TCGA-RCC, randomly selecting proportions for training, validation, and testing. For TCGA-BRCA, we adapted the training and testing slide ID from MSCPT [18]. The detailed description is included in the appendix section.
@@
ABMIL [21] 461K 69.24$\pm$3.90 61.72$\pm$3.36 72.77$\pm$3.15
CLAM-SB [34] 660K 67.80$\pm$5.14 60.51$\pm$5.01 72.46$\pm$4.36
CLAM-MB [34] 660K 60.81$\pm$4.87 55.48$\pm$4.96 67.31$\pm$4.19
TransMIL [52] 2.54M 65.62$\pm$3.20 60.75$\pm$4.04 67.52$\pm$4.16
DSMIL [26] 462K 66.18$\pm$3.08 59.35$\pm$3.18 67.52$\pm$1.56
RRT-MIL [59] 2.63M 66.33$\pm$4.30 61.14$\pm$5.93 71.21$\pm$8.94
CoOp [75] 337K 68.86$\pm$4.35 61.64$\pm$2.40 71.08$\pm$3.22
CoCoOp [74] 370K 69.13$\pm$4.27 61.48$\pm$2.62 72.41$\pm$1.87
Metaprompt [72] 360K 69.12$\pm$4.46 63.39$\pm$4.38 74.65$\pm$7.20
TOP [46] 2.11M 69.74$\pm$3.14 63.39$\pm$4.62 74.41$\pm$5.27
ViLa-MIL [53] 2.77M 72.25$\pm$6.16 62.04$\pm$2.38 75.01$\pm$6.14
MSCPT [18] 1.35M 74.56$\pm$4.54 65.59$\pm$1.85 75.82$\pm$2.38
@@
ABMIL [21] 461K 72.41$\pm$4.25 63.04$\pm$3.62 74.09$\pm$4.38
CLAM-SB [34] 660K 72.34$\pm$6.17 65.51$\pm$3.28 76.16$\pm$4.36
CLAM-MB [34] 660K 73.41$\pm$3.76 66.11$\pm$1.94 77.88$\pm$2.30
TransMIL [52] 2.54M 74.98$\pm$6.01 67.50$\pm$6.00 77.04$\pm$6.14
DSMIL [26] 462K 71.44$\pm$2.72 64.48$\pm$1.64 75.26$\pm$2.28
RRT-MIL [59] 2.63M 71.21$\pm$6.46 64.15$\pm$1.38 75.92$\pm$5.10
CoOp [75] 337K 71.53$\pm$2.45 64.84$\pm$2.40 74.22$\pm$5.02
CoCoOp [74] 370K 72.65$\pm$4.63 66.63$\pm$3.55 66.98$\pm$3.35
Metaprompt [72] 360K 74.86$\pm$4.25 65.03$\pm$1.81 77.88$\pm$3.22
TOP [46] 2.11M 76.13$\pm$6.01 66.55$\pm$1.72 78.58$\pm$5.30
ViLa-MIL [53] 2.77M 74.06$\pm$4.62 66.03$\pm$1.81 78.12$\pm$4.88
MSCPT [18] 1.35M 75.55$\pm$5.25 67.46$\pm$2.43 79.14$\pm$2.63
@@ 4.2 Comparison to State-of-the-Art
, ABMIL [21], CLAM [34], TransMIL [52], DSMIL [26], GTMIL [73], DTMIL [70], RRT-MIL [59] and IBMIL [31], and vision-language methods, including CoOp [75], CoCoOp [74], Metaprompt [72], TOP [46], ViLa-MIL [53], MSCPT [18], QUILT [20], CONCH [32]. Among these, QUILT and CONCH are foundation VLMs.
@@
We validated our whole-slide vision foundation model, pre-trained on 1.3 billion pathology images, using the PLIP text encoder. By incorporating pathology-specific features from Prov-GigaPath [66], the integrated MGPath (PLIP-G) model demonstrated strong performance across multiple metrics on the TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA datasets. When compared to other foundation VLMs such as CONCH and QUILT, our model consistently outperforms them. For example, we achieve a 3% improvement in AUC over CONCH on both the TCGA-BRCA and TCGA-NSCLC datasets.
@@
ABMIL [21] 461K 60.5$\pm$15.9 56.8$\pm$11.8 61.2$\pm$6.1 83.6$\pm$3.1 64.4$\pm$4.2 65.7$\pm$4.7
CLAM-SB [34] 660K 66.7$\pm$13.6 59.9$\pm$13.8 64.0$\pm$7.7 90.1$\pm$2.2 75.3$\pm$7.4 77.6$\pm$7.0
CLAM-MB [34] 660K 68.8$\pm$12.5 60.3$\pm$11.1 63.0$\pm$9.3 90.9$\pm$4.1 76.2$\pm$4.4 78.6$\pm$4.9
TransMIL [52] 2.54M 64.2$\pm$8.5 57.5$\pm$6.4 59.7$\pm$5.4 89.4$\pm$5.6 73.0$\pm$7.8 75.3$\pm$7.2
DSMIL [26] 462K 67.9$\pm$8.0 61.0$\pm$7.0 61.3$\pm$7.0 87.6$\pm$4.5 71.5$\pm$6.6 72.8$\pm$6.4
GTMIL [73] N/A 66.0$\pm$15.3 61.1$\pm$12.3 63.8$\pm$9.9 81.1$\pm$13.3 71.1$\pm$15.7 76.1$\pm$12.9
DTMIL [70] 986.7K 67.5$\pm$10.3 57.3$\pm$11.3 66.6$\pm$7.5 90.0$\pm$4.6 74.4$\pm$5.3 76.8$\pm$5.2
IBMIL [31] N/A 69.2$\pm$7.4 57.4$\pm$8.3 66.9$\pm$6.5 90.5$\pm$4.1 75.1$\pm$5.2 77.2$\pm$4.2
ViLa-MIL [53] 8.8M/47M 74.7$\pm$3.5 67.0$\pm$4.9 67.7$\pm$4.4 92.6$\pm$3.0 78.3$\pm$6.9 80.3$\pm$6.2
CONCH ([32]) 110M 89.46$\pm$10.2 78.5$\pm$9.31 78.78$\pm$9.1 88.08$\pm$4.59 78.21$\pm$4.2 71.67$\pm$19.4
QUILT [20] 63M 79.66$\pm$13.19 72.30$\pm$13.35 72.42$\pm$13.24 96.92$\pm$1.6 78.46$\pm$5.55 86.34$\pm$1.56
@@ 4.4 Ablation Studies
We validate the use of Prov-GigaPath and PLIP under the following settings: (i) using full vision-language PLIP model; (ii) combining Prov-GigaPath with PLIP through the MLP layer which was pre-trained on the large-scale dataset; (iii) integrating Prov-GigaPath with PLIP through adaptor layers which were randomly initialized; (iv) utilizing Prov-GigaPath and an adaptor layer to map to the class output and only train MLP and last FFN layer of slide encoder; (v) only using Prov-GigaPath and an MLP layer to map to the class output and train MLP, the query matrix of last layer and the last FFN layer of slide encoder. Table LABEL:tab:gigapath shows that using Prov-GigaPath combined with PLIP boosts the final performance compared to only using PLIP or Prov-GigaPath.
@@ 4.5 Discussion
While we demonstrate significant improvements in few-shot and zero-shot WSI classification across several settings, this paper does not explore other important challenges. For example, how can we scale the current attention mechanism to handle even larger image patches (e.g., using Flash Attention [11]), or extend the model from classification to tumor segmentation tasks [23]. Additionally, the potential for extending GiGaPath to integrate with other large-scale VLM models, such as CONCH [32], remains unexplored.
@@ Appendix A Description of Dataset Splitting
TCGA-BRCA. This dataset contains 1056 whole slide images of breast invasive carcinoma. To conduct fair experiments, we adapted training and testing slides provided by the GitHub repository of MSCPT [18]. In the MSCPT setup, 20% of the dataset was allocated for training, while the remaining 80% (833 slides) served as the test set. A fixed set of 16-shot WSIs was randomly sampled from the training set. Additionally, MSCPT specified the exact training and testing slides used in its experiments. However, there are 35 slides in which we got errors in the pre-processing steps; thus, we replaced those slides with the other ones (same number of WSI per class) downloaded from Cancer Genome Atlas (TCGA) Data Portal (GDC) [60].
TCGA-RCC &\&& TCGA-NSCLC. We adopt the same data splitting as in ViLa-MIL [53], using 16-shot samples for training in each dataset. For testing, 192 samples were used for TCGA-RCC and 197 samples were used for TCGA-NSCLC.

### v2

@@ 3.3.1 Patches-based Prompting
The frozen image encoder $E_{I}(.)$ (or $A_{I}(E_{I}(.))$ in case of GigaPath-PLIP) is used to map patches $I$ into a feature vector $H=\{H^{(l)}\in\mathbb{R}^{N_{l}\times d},H^{(h)}\in\mathbb{R}^{N_{h}\times d}\}$ where $d$ denotes the feature dimension. To effectively consolidate the extensive set of patch features into a final slide-level representation, we introduce a set of learnable visual prompts $\mathbf{p}_{v}\in\mathbb{R}^{N_{p}\times d}$, which facilitate the progressive merging of patch features in $H^{(l)}$(similarly for $H^{(h)}$) (Figure 2). In particular, we formulate $\mathbf{p}_{v}$ as Query and take all features in $H^{(l)}$ as the Keys $K^{(l)}_{p}$ and Values $V^{(l)}_{p}$ in in self-attention Vaswani (2017). We then associate $\mathbf{p}_{v}$ with patch features as:
@@ 3.3.2 Spatial Patch Group-based Prompting
We next design a trainable message-passing network $g_{\epsilon}(.)$ based on the graph attention layer (GAT) Veličković et al. (2017) to capture the feature representation of each node and its local neighbors. The message passing of the GAT layer is formulated as:
@@ 3.4 Optimal Transport for Visual-Text Alignment
Given descriptive text prompts $\mathbf{T}^{(l)}$ and $\mathbf{T}^{(h)}$ (Eq.(2)) and visual prompt-guided slide features $\mathbf{p}_{v}^{(l)}$ and $\mathbf{p}_{v}^{(h)}$ (Eq.(6)) for low and high resolutions, our goal is to maximize the similarity between slide and text embeddings for each class $c$. Rather than relying on cosine distance, as in prior works Zhou et al. (2022b, a); Zhao et al. (2024); Qu et al. (2024a, a); Singh and Jaggi (2020), we propose using optimal transport (OT)-based distance to capture a more nuanced cross-alignment between visual and text domains. Although OT has been explored for prompt learning in natural images and multi-modal learning Kim et al. (2023); Chen et al. (2023); Nguyen et al. (2024a); Séjourné et al. (2023), we are the first to adapt it for whole-slide imaging (WSI), effectively handling the alignment of multi-magnification patches to capture rich structural details across scales.
@@ 4.1 Settings
PatchGastricADC22Tsuneki and Kanavati (2022) consists of approximately 262K patches derived from WSI of H&E-stained gastric adenocarcinoma specimens, each paired with associated diagnostic captions collected from the University of Health and Welfare, Mita Hospital, Japan. QUILT-1M Ikezogwo et al. (2024) includes approximately 653K images and one million pathology image-text pairs, gathered from 1,087 hours of educational histopathology videos presented by pathologists on YouTube. ARCH Gamper and Rajpoot (2021) is a pathology multiple-instance captioning dataset containing pathology images at the bag and tile level. However, our work focuses on tile-level images from all datasets for our contrastive training strategy. In total, we collected approximately 923K images from these datasets.
@@
For the classification task, the proposed method was evaluated in three datasets from the Cancer Genome Atlas Data PortalThe Cancer Genome Atlas () (TCGA): TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA. We followed the ViLa-MILShi et al. (2024) experimental settings for TCGA-NSCLC and TCGA-RCC, randomly selecting proportions for training, validation, and testing. For TCGA-BRCA, we adapted the training and testing slide ID from MSCPT Han et al. (2024). The detailed description is included in the appendix section.
@@
ABMIL Ilse et al. (2018) 461K 69.24$\pm$3.90 61.72$\pm$3.36 72.77$\pm$3.15
CLAM-SB Lu et al. (2021) 660K 67.80$\pm$5.14 60.51$\pm$5.01 72.46$\pm$4.36
CLAM-MB Lu et al. (2021) 660K 60.81$\pm$4.87 55.48$\pm$4.96 67.31$\pm$4.19
TransMIL Shao et al. (2021) 2.54M 65.62$\pm$3.20 60.75$\pm$4.04 67.52$\pm$4.16
DSMIL Li et al. (2021) 462K 66.18$\pm$3.08 59.35$\pm$3.18 67.52$\pm$1.56
RRT-MIL Tang et al. (2024) 2.63M 66.33$\pm$4.30 61.14$\pm$5.93 71.21$\pm$8.94
CoOp Zhou et al. (2022b) 337K 68.86$\pm$4.35 61.64$\pm$2.40 71.08$\pm$3.22
CoCoOp Zhou et al. (2022a) 370K 69.13$\pm$4.27 61.48$\pm$2.62 72.41$\pm$1.87
Metaprompt Zhao et al. (2024) 360K 69.12$\pm$4.46 63.39$\pm$4.38 74.65$\pm$7.20
TOP Qu et al. (2024a) 2.11M 69.74$\pm$3.14 63.39$\pm$4.62 74.41$\pm$5.27
ViLa-MIL Shi et al. (2024) 2.77M 72.25$\pm$6.16 62.04$\pm$2.38 75.01$\pm$6.14
MSCPT Han et al. (2024) 1.35M 74.56$\pm$4.54 65.59$\pm$1.85 75.82$\pm$2.38
@@
ABMIL Ilse et al. (2018) 461K 72.41$\pm$4.25 63.04$\pm$3.62 74.09$\pm$4.38
CLAM-SB Lu et al. (2021) 660K 72.34$\pm$6.17 65.51$\pm$3.28 76.16$\pm$4.36
CLAM-MB Lu et al. (2021) 660K 73.41$\pm$3.76 66.11$\pm$1.94 77.88$\pm$2.30
TransMIL Shao et al. (2021) 2.54M 74.98$\pm$6.01 67.50$\pm$6.00 77.04$\pm$6.14
DSMIL Li et al. (2021) 462K 71.44$\pm$2.72 64.48$\pm$1.64 75.26$\pm$2.28
RRT-MIL Tang et al. (2024) 2.63M 71.21$\pm$6.46 64.15$\pm$1.38 75.92$\pm$5.10
CoOp Zhou et al. (2022b) 337K 71.53$\pm$2.45 64.84$\pm$2.40 74.22$\pm$5.02
CoCoOp Zhou et al. (2022a) 370K 72.65$\pm$4.63 66.63$\pm$3.55 66.98$\pm$3.35
Metaprompt Zhao et al. (2024) 360K 74.86$\pm$4.25 65.03$\pm$1.81 77.88$\pm$3.22
TOP Qu et al. (2024a) 2.11M 76.13$\pm$6.01 66.55$\pm$1.72 78.58$\pm$5.30
ViLa-MIL Shi et al. (2024) 2.77M 74.06$\pm$4.62 66.03$\pm$1.81 78.12$\pm$4.88
MSCPT Han et al. (2024) 1.35M 75.55$\pm$5.25 67.46$\pm$2.43 79.14$\pm$2.63
@@ 4.2 Comparison to State-of-the-Art
, ABMIL Ilse et al. (2018), CLAM Lu et al. (2021), TransMIL Shao et al. (2021), DSMIL Li et al. (2021), GTMIL Zheng et al. (2022), DTMIL Zhang et al. (2022a), RRT-MIL Tang et al. (2024) and IBMIL Lin et al. (2023), and vision-language methods, including CoOp Zhou et al. (2022b), CoCoOp Zhou et al. (2022a), Metaprompt Zhao et al. (2024), TOP Qu et al. (2024a), ViLa-MIL Shi et al. (2024), MSCPT Han et al. (2024), QUILT Ikezogwo et al. (2024), CONCH Lu et al. (2024). Among these, QUILT and CONCH are foundation VLMs.
@@
We validated our whole-slide vision foundation model, pre-trained on 1.3 billion pathology images, using the PLIP text encoder. By incorporating pathology-specific features from Prov-GigaPath Xu et al. (2024), the integrated MGPath (PLIP-G) model demonstrated strong performance across multiple metrics on the TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA datasets. When compared to other foundation VLMs such as CONCH and QUILT, our model consistently outperforms them. For example, we achieve a 3% improvement in AUC over CONCH on both the TCGA-BRCA and TCGA-NSCLC datasets.
@@
ABMIL Ilse et al. (2018) 461K 60.5$\pm$15.9 56.8$\pm$11.8 61.2$\pm$6.1 83.6$\pm$3.1 64.4$\pm$4.2 65.7$\pm$4.7
CLAM-SB Lu et al. (2021) 660K 66.7$\pm$13.6 59.9$\pm$13.8 64.0$\pm$7.7 90.1$\pm$2.2 75.3$\pm$7.4 77.6$\pm$7.0
CLAM-MB Lu et al. (2021) 660K 68.8$\pm$12.5 60.3$\pm$11.1 63.0$\pm$9.3 90.9$\pm$4.1 76.2$\pm$4.4 78.6$\pm$4.9
TransMIL Shao et al. (2021) 2.54M 64.2$\pm$8.5 57.5$\pm$6.4 59.7$\pm$5.4 89.4$\pm$5.6 73.0$\pm$7.8 75.3$\pm$7.2
DSMIL Li et al. (2021) 462K 67.9$\pm$8.0 61.0$\pm$7.0 61.3$\pm$7.0 87.6$\pm$4.5 71.5$\pm$6.6 72.8$\pm$6.4
GTMIL Zheng et al. (2022) N/A 66.0$\pm$15.3 61.1$\pm$12.3 63.8$\pm$9.9 81.1$\pm$13.3 71.1$\pm$15.7 76.1$\pm$12.9
DTMIL Zhang et al. (2022a) 986.7K 67.5$\pm$10.3 57.3$\pm$11.3 66.6$\pm$7.5 90.0$\pm$4.6 74.4$\pm$5.3 76.8$\pm$5.2
IBMIL Lin et al. (2023) N/A 69.2$\pm$7.4 57.4$\pm$8.3 66.9$\pm$6.5 90.5$\pm$4.1 75.1$\pm$5.2 77.2$\pm$4.2
ViLa-MIL Shi et al. (2024) 8.8M/47M 74.7$\pm$3.5 67.0$\pm$4.9 67.7$\pm$4.4 92.6$\pm$3.0 78.3$\pm$6.9 80.3$\pm$6.2
CONCH (Lu et al. (2024)) 110M 89.46$\pm$10.2 78.5$\pm$9.31 78.78$\pm$9.1 88.08$\pm$4.59 78.21$\pm$4.2 71.67$\pm$19.4
QUILT Ikezogwo et al. (2024) 63M 79.66$\pm$13.19 72.30$\pm$13.35 72.42$\pm$13.24 96.92$\pm$1.6 78.46$\pm$5.55 86.34$\pm$1.56
@@ 4.4 Ablation Studies
We validate the use of Prov-GigaPath and PLIP under the following settings: (i) using full vision-language PLIP model; (ii) combining Prov-GigaPath with PLIP through the MLP layer which was pre-trained on the large-scale dataset; (iii) integrating Prov-GigaPath with PLIP through adaptor layers which were randomly initialized; (iv) utilizing Prov-GigaPath and an adaptor layer to map to the class output and only train MLP and last FFN layer of slide encoder; (v) only using Prov-GigaPath and an MLP layer to map to the class output and train MLP, the query matrix of last layer and the last FFN layer of slide encoder. Table 5 shows that using Prov-GigaPath combined with PLIP boosts the final performance compared to only using PLIP or Prov-GigaPath.
@@ 4.5 Discussion
While we demonstrate significant improvements in few-shot and zero-shot WSI classification across several settings, this paper does not explore other important challenges. For example, how can we scale the current attention mechanism to handle even larger image patches (e.g., using Flash Attention Dao et al. (2022)), or extend the model from classification to tumor segmentation tasks Khened et al. (2021). Additionally, the potential for extending GiGaPath to integrate with other large-scale VLM models, such as CONCH Lu et al. (2024), remains unexplored.
@@ 5 Conclusion
6 Acknowledgement
This work is supported in part by funds from the German Ministry of Education and Research (BMBF) under grant agreements No. 01D2208A and No. 01KD2414A (project FAIrPaCT).
The project “Development of an intelligent collaboration service for AI-based collaboration between rescue services and central emergency rooms” (acronym: CONNECT_ED) is funded by the German Federal Ministry of Education and Research under grant number 16SV8977 and by the joint project KISSKI under grant number 1IS22093E.
The authors gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy/Grete at NHR-Nord@Göttingen as part of the NHR infrastructure. The calculations for this research were conducted with computing resources under the project nim00014.
The authors gratefully acknowledge the computing time granted by the KISSKI project. The calculations for this research were conducted with computing resources under the project kisski-umg-fairpact-2.
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Duy M. H. Nguyen. Duy M. H. Nguyen and Daniel Sonntag are also supported by the XAINES project (BMBF, 01IW20005), No-IDLE project (BMBF, 01IW23002), and the Endowed Chair of Applied Artificial Intelligence, Oldenburg University.
@@ Appendix A Description of Dataset Splitting
TCGA-BRCA. This dataset contains 1056 whole slide images of breast invasive carcinoma. To conduct fair experiments, we adapted training and testing slides provided by the GitHub repository of MSCPT Han et al. (2024). In the MSCPT setup, 20% of the dataset was allocated for training, while the remaining 80% (833 slides) served as the test set. A fixed set of 16-shot WSIs was randomly sampled from the training set. Additionally, MSCPT specified the exact training and testing slides used in its experiments. However, there are 35 slides in which we got errors in the pre-processing steps; thus, we replaced those slides with the other ones (same number of WSI per class) downloaded from Cancer Genome Atlas (TCGA) Data Portal (GDC) The Cancer Genome Atlas () (TCGA).
TCGA-RCC &\&& TCGA-NSCLC. We adopt the same data splitting as in ViLa-MIL Shi et al. (2024), using 16-shot samples for training in each dataset. For testing, 192 samples were used for TCGA-RCC and 197 samples were used for TCGA-NSCLC.